OpenStudy (anonymous):

1.   Find the rational roots of x 4 + 8x 3 + 7x 2 –40x–60 = 0.

OpenStudy (amtran_bus):

I would begin by factoring.

OpenStudy (anonymous):

try all numbers that divide \(60\) i would try \(-1\) first

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

unfortunately \(-1\) does not work try \(-2\) instead, that does work

OpenStudy (anonymous):

{±factors of last term/factors of first term} {±1,2,3,4,5,6,10,12,15,20,30,60/1} Test each root to see which one works. Try -6 and apply synthetic division. 1...8...7...-40...-60 ....-6.-12..30....60 1...2..-5...-10....0 Rewrite the remaining expression. x³ + 2x² - 5x - 10 = 0 Factor. x²(x + 2) - 5(x + 2) = 0 Combine. (x² - 5)(x + 2) = 0 Set each by zero and solve for x. x² - 5 = 0 and x + 2 = 0 You get: x = {±√5, -2} However, ±√5 are not rational because it does not express in fraction. Therefore.... The rational roots are -6 and -2. Hope that helps!

OpenStudy (anonymous):

also \(-6\) i think and those are the only rational ones

OpenStudy (anonymous):

hope my explanation helps well explanation and answer!